# SPAND: Sleep Prediction Architecture using Network Dynamics

### 20 Jan 2024 v1

@@ Abstract.
Sleep behavior significantly impacts health and acts as an indicator of physical and mental well-being. Monitoring and predicting sleep behavior with ubiquitous sensors may therefore assist in both sleep management and tracking of related health conditions. While sleep behavior depends on, and is reflected in the physiology of a person, it is also impacted by external factors such as digital media usage, social network contagion, and the surrounding weather. In this work, we propose SPAND (Sleep Prediction Architecture using Network Dynamics), a system that exploits social contagion in sleep behavior through graph networks and integrates it with physiological and phone data extracted from ubiquitous mobile and wearable devices for predicting next-day sleep labels about sleep duration. Our architecture overcomes the limitations of large-scale graphs containing connections irrelevant to sleep behavior by devising an attention mechanism. The extensive experimental evaluation highlights the improvement provided by incorporating social networks in the model. Additionally, we conduct robustness analysis to demonstrate the system’s performance in real-life conditions. The outcomes affirm the stability of SPAND against perturbations in input data. Further analyses emphasize the significance of network topology in prediction performance revealing that users with higher eigenvalue centrality are more vulnerable to data perturbations.
@@
ccs: Human-centered computing mobile computingccs: Applied computing Health care information systems.
@@ 1. Introduction
Sleep behavior is closely associated with physical(Reid et al., 2006)(Gothe et al., 2020) and mental health(Hosker et al., 2019)(João et al., 2018), both as an influence(Cappuccio and Miller, 2017) and as an effect(Morris et al., 2015). Short sleep duration can increase the risk of multiple diseases and worsen existing conditions such as cardio-metabolic dysfunction(Cappuccio and Miller, 2017), obesity(Emsellem et al., 2014), hormonal imbalance, and diabetes(Schmid et al., 2007) and impact a person’s functioning in daily life including impaired work efficiency(van Oostveen and Vermeulen, 2014)(Krueger, 1989), fatigue, sleepiness, and performance deficits(Kryger et al., 2010).
@@
To address these issues, mobile health (mHealth) applications are proposed for monitoring and providing appropriate interventions. The interventions that utilize behavior change techniques such as shaping knowledge, goals, and planning have been shown to have desirable effects(Arroyo and Zawadzki, 2022) on sleep behavior and have great potential to mitigate the above-mentioned risks. These mHealth systems require accurate monitoring and prediction of sleep behavior. Sleep duration, when used in machine learning models as an input feature, can improve well-being prediction(Taylor et al., 2017). The prediction of sleep duration itself, however, can be a challenging task. While time-of-day and recent sleep/wake timing are strong predictors for the next sleep episode duration, several external factors, such as social interactions and digital media usage, can also impact sleep duration (Hale et al., 2019)(Orzech et al., 2016). The use of digital media can impact sleep duration directly or can lead to a person changing their behavior based on the people in their social network. There is a complex dynamic between an individual’s social network, their digital media interactions encompassing calls, SMS messages, and app usage, and the influence these factors exert on that individual’s behaviors.
@@
Previous studies have established the existence of sleep behavior ”contagion” in a social network. Using a network of 8349 adolescents to investigate the relationship between sleep behavior and drug usage(Mednick et al., 2010), Mednick and colleagues identified clusters of people with poor sleep behavior and demonstrated, through statistical analysis, how network centrality impacts sleep behavior outcomes. Additionally, these investigators found a connection between a person’s future sleep duration and their friend’s current sleep duration(Mednick et al., 2010). Another aspect of the environment is weather which has also been shown to affect sleep duration (Ferguson et al., 2023)(O’Connell et al., 2014)(Bentley et al., 2013). Weather impacts physical activity and sedentary behavior(Tucker and Gilliland, 2007)(Turrisi et al., 2021), which in turn impacts sleep behavior(Mendelson et al., 2016)(Murray et al., 2017). Weather can also impact well-being and mood(Denissen et al., 2008), both of which can affect sleep behavior and is, therefore, an important variable to consider when predicting sleep duration.
@@
To this end, we use Graph Convolution Networks (GCN) to aggregate features from the user’s neighbors and predict sleep duration labels for the entire network. However, irrelevant nodes (e.g., nodes that do not contribute to a user’s sleep duration behavior) or noisy graphs can hinder prediction performance. To alleviate this problem, we incorporate an additional attention mechanism that filters out irrelevant nodes by paying more ”attention” to important nodes by utilizing Graph Attention Networks (GAN). Another limitation of graph-based architecture is that it requires both a feature matrix and a graph network as inputs; this limits the size of graphs that can be handled by the model. However, in real life, the magnitude of social interactions often changes. To address this limitation, we utilize GEDD: Graph Extraction for Dynamic Graphs, an iterative algorithm based on the principle of graph convolution networks to transform a set of varying size graphs into a set of graphs with fixed predetermined size without any loss of the user or graph information. Lastly, we extract the temporal trends in the data through a Long Short-term Memory (LSTM) network module and integrate both multi-user aggregated features with extracted temporal trends in a deep learning architecture, abbreviated as SPAND, that predicts the next-day sleep duration labels for all the users present in the network.
@@
  • We present SPAND: Sleep Prediction Architecture using Network Dynamics, that exploits contagion in sleep behavior for next-day sleep duration prediction. It utilizes rich multimodal data from wearables, mobile phones, surveys, and weather and integrates social networks in the prediction model. Furthermore, the architecture overcomes the limitations of noisy/missing-user graphs and large networks by leveraging an attention mechanism.
@@ 2.1. Social contagion
Contagion is ”rapid communication of an influence (such as a doctrine or emotional state)”(con, 2023). Contagion in behavior is the production of similar behavior in response to interaction with another person(Hatfield et al., 1992)(Nickerson, 2021)(Centola, 2018). One example is the tendency to mimic and synchronize expressions, vocalizations, postures, and movements with those of another person. Multiple works have explored how health outcomes such as obesity(Christakis and Fowler, 2007), tuberculosis (Klovdahl et al., 2001), pneumonia(Meyers et al., 2003), and severe acute respiratory syndrome(Meyers et al., 2005), can spread in a network suggesting the existence of contagion in health behaviors(Smith and Christakis, 2008). The work in (Mednick et al., 2010) suggests the contagion of sleep behavior in a social network. An empirical study was conducted using data collected from students in grades 7–12 in the National Longitudinal Study of Adolescent Health (Udry, 2003) over 8 years. The dataset was used to extract a social network of 8349 adolescents to study the relationship between sleep behavior and drug usage. The study found clusters of poor sleep behavior; statistical analysis revealed that the probability of a person sleeping less than seven hours goes up by $11\%$ if their friend also sleeps less than seven hours. Further findings indicated that high network centrality negatively impacts future sleep behavior. We extend this work to predict next-day sleep duration labels at a day-to-day resolution.
@@ 2.2.1. Wearable technologies
The gold standard to measure sleep is polysomnography(Rundo and Downey III, 2019) which collects multiple physiological parameters, two of which are the electroencephalogram (EEG) and electrocardiogram (ECG). However, it is expensive and burdensome to participants and staff. A large body of literature investigates learning sleep characteristics from the information-rich brain and heart activity signals collected using EEG(Tsinalis et al., 2016)(Liu et al., 2017) and ECG(Wei et al., 2018). However, the collection of EEG and ECG data requires wearing electrodes at multiple locations of the body which is not convenient in real-life settings.
As an alternative to EEG and ECG data, wearable devices are easy to use, less obtrusive, portable, low-cost, and can be worn without any difficulty when the individual chooses to go to bed/sleep. The efficacy of using actigraphy sensors to predict sleep quality was investigated in (Sathyanarayana et al., 2016). Learning sleep attributes from smartwatches is also thoroughly investigated in previous works (Chang et al., 2018)(Fellger et al., 2020). Multi-modal sensing(Zhai et al., 2020) extends this by combining data from the actigraphy sensor and ECG for understanding the time spent in different sleep stages. Our system is similar to most of these works, as we also utilize wearable acceleration sensors for predicting sleep duration. However, in addition to acceleration data, we incorporate the impact of skin temperature, skin conductance, surroundings, and digital media usage on sleep duration by utilizing weather and mobile metadata.
@@ 2.2.2. Non-wearable low user-burden technologies
Another interesting body of literature focuses on solutions in which the user does not have to wear any sensor at all. For instance, a system that uses radio-signal transmitting devices in the room to monitor sleep biomarkers is non-invasive and low-burden (Hsu et al., 2017)(Rahman et al., 2015). Another approach involves multiple sensors throughout a smart home to forecast sleep behavior(Williams and Cook, 2017). However, these solutions are limited to controlled environments, can be costly, and are not effective when there are multiple people in the home or when the participant is not at home (e.g., traveling).
These limitations were overcome by work (Wahl and Amft, 2018)(Wang et al., 2014)(Min et al., 2014) in which only mobile phone data were utilized to learn sleep behaviors. However, these require the smartphone to be always placed close to the user so that features representing mobility, light, and activity can be extracted. Another alternative technology is a bed sensor e.g. EMFIT QS sleep analyzer, however, the current state of this technology does not provide performance comparable to wrist actigraphy(Buimer et al., 2023).
In summary, there are two main types of sensing technologies based on proximity to the user. Close proximity sensors such as smartwatches and mobile phones are used by many people and are relatively low-cost. Distant proximity sensors such as radio transmitters and IoT (Internet-of-Things) devices require specialized setup. Thus, there is a trade-off between user burden and sensor burden (e.g., cost, ubiquity) when choosing the sensing technology. In future work, insights from (Wahl and Amft, 2018)(Hsu et al., 2017) and (Abdullah et al., 2014) may be useful in reducing the burden of wearing the device.
@@ 2.3.2. Estimation using machine learning models
Linear regression models using smartphone-derived features related to mobility, light exposure, and usage were applied to infer sleep duration (Chen et al., 2013)(Lane et al., 2012). EEG and ECG data have been used to identify sleep disorders and classify sleep stages using various models, including k-nearest neighbor classifiers, convolutional neural networks (CNN)(Phan et al., 2018)(Tsinalis et al., 2016)(Liu et al., 2017), deep neural networks(Wei et al., 2018), and support vector machines(Bsoul et al., 2010). However, these models often relied on handcrafted features and may have limited generalizability. Another work combined (Zhai et al., 2020) actigraphy data with ECG data for calculating the duration of time spent in different sleep stages using multiple machine learning models.
@@ 3. Data Collection
Gender Age
Mean age = 21 years with
standard deviation= 1.3 years
Male 158 17-22 240
Female 92 22-28 10
Table 1. Demographics Statistics
@@
The dataset was collected from 250 participants (total  7,500 days of data) from 2013 to 2017. Participants were asked to identify 5 people whom they usually call or SMS at least once a week and who meet the following criteria (i) undergraduate student (ii) own an Android smartphone (iii) age between 18 and 60 years. People could not participate if they met any of the following conditions (i) people who have a problem wearing wrist sensors e.g., irritated skin on the wrist, etc (ii) pregnant women (iii) freshmen (cohort 1 and 3 only) (iv) people who traveled more than one time zone away one week prior to the study or have plans to travel more than once zone away during the protocol (except for cohort 7). Some participants (not exhaustively) need to be connected through call or SMS at least once a week. For the last two cohorts, their friends who were not enrolled in the main study also joined as friends of participants even though they were not Android users. We preferred recruiting bigger groups of people who were friends with each other. Each person participated for  30 consecutive days (cohorts 1-6) or  110 consecutive days (cohort 7) during one academic term.; cohorts 1-6 had more participants that participated for a shorter duration, while cohort 7 had the least number of participants who were studied the longest. All participants in the study were students at one university. The demographic statistics are listed in Table.1. Each academic term is considered a cohort and cohort number is assigned in chronological order i.e. cohort 1 was the first one and cohort 7 was the last one in the study. The number of participants in each cohort from 1-7 was 20, 48, 46, 47, 40, 35, and 15, respectively. There were available data for 15-25 days from most participants in cohorts 1-6 and for 40-60 days from most participants in the last cohort.
@@
During the study period, the participants (i) wore a wrist device that collected skin conductance, skin temperature, and acceleration (Q-sensor, Affective, USA) on their dominant hand; (ii) had an app (funf(fun, 2011)) installed on their phone that collected metadata of calls, SMS, screen usage, and location. During these years of study, there were very few phone apps on which SMS or calls could be made, and therefore, our data collection about texting and calls made using the phone was considered to be complete; (iii) answered morning and evening daily surveys about drugs and alcohol intake, sleep time, naps, exercise, and academic and extracurricular activities. At the beginning and the end of the study, participants completed social network surveys including questions about social activity partners, emergency contacts, and roommates to identify people that participants would often interact with. For more details about the data collected in the study, please refer to Supplementary Material section D.
The study protocols and informed consent procedure were approved by [institutions (will replace after acceptance for anonymous review)]. All participants signed an informed consent form. All methods were performed in accordance with the relevant guidelines and regulations complying with the declaration of Helsinki. To protect study participants’ privacy and consent and since some of the participants did not consent to sharing their data with the third-party researchers, raw data will not be publicly available. However, deidentified data including the processed input features, labels, and graph networks extracted from phone data, which are used to evaluate the models in this work and would be needed for reproducibility, are available online(Khalid and Sano, 2023b).
@@
The final dataset was deidentified and each participant was assigned a unique ID. Windowed features extracted from wearable device data contained multiple features for activity and physiology. Phone calls, SMS, and screen time were also summed over windows and used as features (e.g., number of events, timing, number of people communicated, distance traveled within a day or within different timeframes). Multiple features explaining weather conditions (e.g., temperature, precipitation, cloud cover, pressure, humidity, sunlight, and wind) were added to the feature matrix. After removing all sleep-behavior labels, such as sleep efficiency, data from all modalities were concatenated to create one feature matrix with 317 features (details in Supplementary Material D.1.1). This dataset is available at (Khalid and Sano, 2023b).
4. SPAND: Sleep Prediction Architecture using Network Dynamics

### 27 Jan 2024 v2

# SleepNet: Attention-Enhanced Robust Sleep Prediction using Dynamic Social Networks

@@ Abstract.
Sleep behavior significantly impacts health and acts as an indicator of physical and mental well-being. Monitoring and predicting sleep behavior with ubiquitous sensors may therefore assist in both sleep management and tracking of related health conditions. While sleep behavior depends on, and is reflected in the physiology of a person, it is also impacted by external factors such as digital media usage, social network contagion, and the surrounding weather. In this work, we propose SleepNet, a system that exploits social contagion in sleep behavior through graph networks and integrates it with physiological and phone data extracted from ubiquitous mobile and wearable devices for predicting next-day sleep labels about sleep duration. Our architecture overcomes the limitations of large-scale graphs containing connections irrelevant to sleep behavior by devising an attention mechanism. The extensive experimental evaluation highlights the improvement provided by incorporating social networks in the model. Additionally, we conduct robustness analysis to demonstrate the system’s performance in real-life conditions. The outcomes affirm the stability of SleepNet against perturbations in input data. Further analyses emphasize the significance of network topology in prediction performance revealing that users with higher eigenvalue centrality are more vulnerable to data perturbations.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Human-centered computing mobile computingccs: Applied computing Health care information systems.
@@ 1. Introduction
Sleep behavior is closely associated with physical  (Reid et al., 2006; Gothe et al., 2020) and mental health (Hosker et al., 2019; João et al., 2018), both as an influence (Cappuccio and Miller, 2017) and as an effect (Morris et al., 2015). Short sleep duration can increase the risk of multiple diseases and worsen existing conditions such as cardio-metabolic dysfunction (Cappuccio and Miller, 2017), obesity (Emsellem et al., 2014), and diabetes (Schmid et al., 2007) and impact a person’s functioning in daily life including impaired work efficiency (van Oostveen and Vermeulen, 2014; Krueger, 1989), fatigue, sleepiness, and performance deficits (Kryger et al., 2010).
@@
To address these issues, mobile health (mHealth) applications are proposed for monitoring and providing appropriate interventions. The interventions that utilize behavior change techniques such as shaping knowledge, goals, and planning have been shown to have desirable effects (Arroyo and Zawadzki, 2022) on sleep behavior and have great potential to mitigate the above-mentioned risks. These mHealth systems require accurate monitoring and prediction of sleep behavior. Sleep duration, when used in machine learning models as an input feature, can improve well-being prediction (Taylor et al., 2017). The prediction of sleep duration itself, however, can be a challenging task. While time-of-day and recent sleep/wake timing are strong predictors for the next sleep episode duration, several external factors, such as social interactions and digital media usage, can also impact sleep duration (Hale et al., 2019; Orzech et al., 2016). The use of digital media can impact sleep duration directly or can lead to a person changing their behavior based on the people in their social network. There is a complex dynamic between an individual’s social network, their digital media interactions encompassing calls, SMS messages, and app usage, and the influence these factors exert on that individual’s behaviors.
@@
Previous studies have established the existence of sleep behavior ”contagion” in a social network. Using a network of 8349 adolescents to investigate the relationship between sleep behavior and drug usage (Mednick et al., 2010), Mednick and colleagues identified clusters of people with poor sleep behavior and demonstrated, through statistical analysis, how network centrality impacts sleep behavior outcomes. Additionally, these investigators found a connection between a person’s future sleep duration and their friend’s current sleep duration (Mednick et al., 2010). Another aspect of the environment is weather which has also been shown to affect sleep duration (Ferguson et al., 2023; O’Connell et al., 2014; Bentley et al., 2013). Weather impacts physical activity and sedentary behavior (Tucker and Gilliland, 2007; Turrisi et al., 2021), which in turn impacts sleep behavior (Mendelson et al., 2016; Murray et al., 2017). Weather can also impact well-being and mood (Denissen et al., 2008), both of which can affect sleep behavior and is, therefore, an important variable to consider when predicting sleep duration.
@@
To this end, we use Graph Convolution Networks (GCN) to aggregate features from the user’s neighbors and predict sleep duration labels for the entire network. However, irrelevant nodes (e.g., nodes that do not contribute to a user’s sleep duration behavior) or noisy graphs can hinder prediction performance. To alleviate this problem, we incorporate an additional attention mechanism that filters out irrelevant nodes by paying more ”attention” to important nodes by utilizing Graph Attention Networks (GAN). Another limitation of graph-based architecture is that it requires both a feature matrix and a graph network as inputs; this limits the size of graphs that can be handled by the model. However, in real life, the magnitude of social interactions often changes. To address this limitation, we utilize GEDD: Graph Extraction for Dynamic Graphs, an iterative algorithm based on the principle of graph convolution networks to transform a set of varying size graphs into a set of graphs with fixed predetermined size without any loss of the user or graph information. Lastly, we extract the temporal trends in the data through a Long Short-term Memory (LSTM) network module and integrate both multi-user aggregated features with extracted temporal trends in a deep learning architecture, abbreviated as SleepNet, that predicts the next-day sleep duration labels for all the users present in the network.
@@
  • We present SleepNet, a framework that exploits contagion in sleep behavior for next-day sleep duration prediction. It utilizes rich multimodal data from wearables, mobile phones, surveys, and weather and integrates social networks in the prediction model. Furthermore, the architecture overcomes the limitations of noisy/missing-user graphs and large networks by leveraging an attention mechanism.
@@ 2.1. Social contagion
Contagion is ”rapid communication of an influence (such as a doctrine or emotional state)” (con, 2023). Contagion in behavior is the production of similar behavior in response to interaction with another person (Hatfield et al., 1992; Nickerson, 2021; Centola, 2018). One example is the tendency to mimic and synchronize expressions, vocalizations, postures, and movements with those of another person. Multiple works have explored how health outcomes such as obesity (Christakis and Fowler, 2007), tuberculosis (Klovdahl et al., 2001), pneumonia (Meyers et al., 2003), and severe acute respiratory syndrome (Meyers et al., 2005), can spread in a network suggesting the existence of contagion in health behaviors (Smith and Christakis, 2008). The work in (Mednick et al., 2010) suggests the contagion of sleep behavior in a social network. An empirical study was conducted using data collected from students in grades 7–12 in the National Longitudinal Study of Adolescent Health (Udry, 2003) over 8 years. A social network of 8349 adolescents was extracted to study the relationship between sleep behavior and drug usage. The study found clusters of poor sleep behavior; statistical analysis revealed that the probability of a person sleeping less than seven hours goes up by $11\%$ if their friend also sleeps less than seven hours. Further findings indicated that high network centrality negatively impacts future sleep behavior. We extend this work to predict next-day sleep duration labels at a day-to-day resolution.
@@ 2.2.1. Wearable Technologies
The gold standard to measure sleep is polysomnography (Rundo and Downey III, 2019) which collects multiple physiological parameters, two of which are the electroencephalogram (EEG) and electrocardiogram (ECG). However, it is expensive and burdensome to participants and staff. A large body of literature investigates learning sleep characteristics from the information-rich brain and heart activity signals collected using EEG (Tsinalis et al., 2016; Liu et al., 2017) and ECG (Wei et al., 2018). However, the collection of EEG and ECG data requires wearing electrodes at multiple locations of the body which is not convenient in real-life settings.
As an alternative to EEG and ECG data, wearable devices are easy to use, less obtrusive, portable, low-cost, and can be worn without any difficulty when the individual chooses to go to bed/sleep. The efficacy of using actigraphy sensors to predict sleep quality was investigated in (Sathyanarayana et al., 2016). Learning sleep attributes from smartwatches is also thoroughly investigated in previous works (Chang et al., 2018; Fellger et al., 2020). Multi-modal sensing (Zhai et al., 2020) extends this by combining data from the actigraphy sensor and ECG for understanding the time spent in different sleep stages. Our system is similar to most of these works, as we also utilize wearable acceleration sensors for predicting sleep duration. However, in addition to acceleration data, we incorporate the impact of skin temperature, skin conductance, surroundings, and digital media usage on sleep duration by utilizing weather and mobile metadata.
@@ 2.2.2. Non-wearable Low User-burden Technologies
Another interesting body of literature focuses on solutions in which the user does not have to wear any sensor at all. For instance, a system that uses radio-signal transmitting devices in the room to monitor sleep biomarkers is non-invasive and low-burden (Hsu et al., 2017; Rahman et al., 2015). Another approach involves multiple sensors throughout a smart home to forecast sleep behavior (Williams and Cook, 2017). However, these solutions are limited to controlled environments, can be costly, and are not effective when there are multiple people in the home or when the participant is not at home (e.g., traveling).
These limitations were overcome by work (Wahl and Amft, 2018; Wang et al., 2014; Min et al., 2014) in which only mobile phone data were utilized to learn sleep behaviors. However, these require the smartphone to be always placed close to the user so that features representing mobility, light, and activity can be extracted. Another alternative technology is a bed sensor e.g. EMFIT QS sleep analyzer, however, the current state of this technology does not provide performance comparable to wrist actigraphy (Buimer et al., 2023).
In summary, there are two main types of sensing technologies based on proximity to the user. Close proximity sensors such as smartwatches and mobile phones are used by many people and are relatively low-cost. Distant proximity sensors such as radio transmitters and IoT (Internet-of-Things) devices require specialized setup. Thus, there is a trade-off between user burden and sensor burden (e.g., cost, ubiquity) when choosing the sensing technology. In future work, insights from (Wahl and Amft, 2018; Hsu et al., 2017; Abdullah et al., 2014) may be useful in reducing the burden of wearing the device.
@@ 2.3.2. Estimation using Machine Learning Models
Linear regression models using smartphone-derived features related to mobility, light exposure, and usage were applied to infer sleep duration (Chen et al., 2013; Lane et al., 2012). EEG and ECG data have been used to identify sleep disorders and classify sleep stages using various models, including k-nearest neighbor classifiers, convolutional neural networks (CNN) (Phan et al., 2018; Tsinalis et al., 2016; Liu et al., 2017), deep neural networks (Wei et al., 2018), and support vector machines (Bsoul et al., 2010). However, these models often relied on handcrafted features and may have limited generalizability.
Another work combined (Zhai et al., 2020) actigraphy data with ECG data for calculating the duration of time spent in different sleep stages using multiple machine learning models.
@@ 3. Data Collection
Over a 5-year period from 2013 to 2017, we collected multi-modal data from 250 participants (158 female, 92 male) in the SNAPSHOT study (Sano et al., 2018) with an average age of 21 years (standard deviation: 1.3 years). The majority, about 240 participants, were aged 17-22, while 10 were between 22-28 years old.
Study Design and Participant Cohorts: Each academic term represents a cohort, numbered chronologically from 1 (first) to 7 (last). Each person participated for  30 consecutive days (cohorts 1-6) or  110 consecutive days (cohort 7) during one academic term. The numbers of participants in each cohort from 1-7 were 20, 48, 46, 47, 40, 35, and 15, respectively. After preprocessing, over 80% of the participants in cohorts 1-6 contributed data for 10-25 days, while in the last cohort, data were available for 40-65 days for around 75% of the participants.
@@
Participant Eligibility Criteria: All participants were students at one university. Participants were asked to identify 5 people whom they usually call or SMS at least once a week and all of them must meet the following criteria, (i) undergraduate students, (ii) own an Android smartphone, and (iii) age between 18 and 60 years. Exclusion criteria encompassed individuals with conditions such as wrist sensor discomfort (e.g., irritated skin), pregnant women, freshmen (cohort 1 and 3 only), and those who had traveled more than one time zone away one week before or had plans for such travel during the study (except for cohort 7). In the last two cohorts, friends of participants, even if not Android users, were included, emphasizing the recruitment of interconnected friend groups. We preferred recruiting bigger groups of people who were friends with each other.
@@
Data Collection and Processing: During the study period, the participants (i) wore a wrist device that collected skin conductance, skin temperature, and acceleration (Q-sensor, Affective, USA) on their dominant hand; (ii) had an app (funf (fun, 2011)) installed on their phone that collected metadata of calls, SMS, screen usage, and location; (iii) answered morning and evening daily surveys about drugs and alcohol intake, sleep time, naps, exercise, and academic and extracurricular activities. During these years of study, there were very few phone apps on which SMS or calls could be made, and therefore, our data collection about texting and calls made using the phone was considered to be complete. At the start and the end of the study, participants completed social network surveys including questions about social activity partners, emergency contacts, and roommates to identify people that participants would often interact with. For more details, please refer to Supplementary Material section E.
@@
Subsequently, windowed features were extracted from wearable device data containing multiple features for activity and physiology. Phone calls, SMS, and screen time were also summed over windows and used as features (e.g., number of events, timing, number of people communicated, distance traveled within a day or within different timeframes). Multiple features explaining weather conditions (e.g., temperature, precipitation, cloud cover, pressure, humidity, sunlight, and wind) were added to the feature matrix. After removing all sleep-behavior labels, such as sleep efficiency, data from all modalities were concatenated to create one feature matrix with 317 features (details in Supplementary Material E.1.1).
Ethics Approval and Data Availability: The study protocols and informed consent procedure were approved by the Massachusetts Institute of Technology and Partners HealthCare Institutional Review Boards. The study was registered on clinicaltrials.gov (NCT02846077). All participants signed an informed consent form. All methods were performed in accordance with the relevant guidelines. To protect study participants’ privacy and consent and since some of the participants did not consent to sharing their data with the third-party researchers, raw data will not be publicly available. However, deidentified data including the processed input features, labels, and graph networks extracted from phone data, which are used to evaluate the models in this work and would be needed for reproducibility, are available online (Khalid and Sano, 2023b).
4. SleepNet: Sleep Prediction Architecture using Network Dynamics